# Detours for Navigating Instructional Videos

### v1

@@ 4.1 Detour video retrieval
We benchmark models on finding the correct detour video given the source video and the query, from amongst all videos in the test set (3,873 videos).
@@
These three variants capture whether individual embeddings (or simple combinations of them) are sufficient for detour retrieval.
@@
Ablations. Similar to how we evaluate the baselines with different inputs, we compare the performance of our method (i.e. “with $V_{s},\mathcal{Q}$”) with ablations — “with $V_{s}$” and “with $\mathcal{Q}$”. These baselines capture the performance when no query is provided and when source video $V_{s}$ is not given to the model, respectively. Next, recall that we use a language model to generate timestamps for both the summaries and the detour timestamps (see Fig. 2 and Sec. 3.2) that can be noisy. Instead, for this ablation, we parse the summaries and queries and assign timestamps based on the similarity score between the narrations and the summaries. The language model does not handle timestamps for this “Ours w/ parser” baseline. Finally, we evaluate the retrieval performance without sampling negatives from the same task i.e. no hard negatives in “Ours w/o hard negatives” baseline.
@@
Metrics. Following standard retrieval evaluation [54, 55, 73, 52], we report recall@$k$ for $k\in[5,10,50]$ and the median rank. Recall@$k$ ranges in $[0,1]$, higher the better. The median rank ranges between $1$ and the size of the retrieval candidate set (i.e. $3,873$), lower the better.
@@
Results. We report the results with both inputs, i.e. “with $V_{s},\mathcal{Q}$” in Table 1 (left) for all baselines on all metrics. We also report medR for different inputs in Table 2 (left). Our method outperforms all the baselines by a significant margin. Our performance gain is 7.9% w.r.t. InternVideo [70] at R@5 and the gap further increases to 35.2% for R@50. Across all the different input combinations, InternVideo [70] “with $\mathcal{Q}$” is the second best and attains a medR of 138 — much lower than our medR of 30 (Tab. 2).
@@
We attribute our large gain to three crucial factors — (a) thanks to its tokens and long sequence length in the LLM, our model is capable of long video understanding, unlike CoVR [67], which samples a few frames from the video, (b) appropriately using the source video context $V_{d}$, unlike CLIP [61], CLIP-Hitchhiker [7], and InternVideo [70] since our model is more than a standard text-to-video retrieval model and (c) conditioning the retrieval on the query text.
@@
Fig. 5 shows example detours inferred by our model. In all the cases, our retrieval model ranks the correct detour video at top-1. It is evident that detour video retrieval is more challenging than conventional text-to-video-retrieval because the detour query can have missing context, e.g. “how do I add garnish after pouring?” does not reveal the video task; the previous video context is needed for the correct retrieval. See Supp. for more visualizations, comparisons and some failure cases. The performance trend is similar for common and novel tasks of the dataset, reinforcing that training on a sufficiently large dataset with many recipes enables good generalization. We show the performance of each test subset in the Supp.
@@ 4.2 Detour window localization
Method
$V_{s}$
$\mathcal{Q}$
MedR $\downarrow$
CLIP [61]
314
@@
Baselines. While there are no existing works that use previous minutes-long viewing history along with the query to perform temporal localization, we use state-of-the-art localization works as our baselines and also strengthen them by providing source video context, as described below.
@@
  • 2D-TAN [83], VSLNet [81], UMT [46]: All these prior models aim to localize text in videos. Thus, we also propose variants of these methods at different input combinations, same as the detour retrieval task. To provide video context, we use a visual mapper $f_{vm}$ similar to our proposed approach—thus providing a late fusion of video $V_{s}$ context. These visual tokens are prepended to the text token, same as our token sequence (see Fig. 4). This enhancement enables us to evaluate “with $V_{s}$” and “with $V_{s}$, $\mathcal{Q}$” in addition to the standard “with $\mathcal{Q}$
@@
Results. Tab. 1 shows the results with maximum supervision (“with $V_{s}$, $\mathcal{Q}$”) and Tab. 2 (right) compares against different input combinations. We use the same set of ablations as the detour retrieval task, see Sec. 4.1. We clearly outperform all the baselines and ablations by a clear margin. Our mean recall@1 is 3.2% more than the second best performing method STALE [57]. The same trend is true for all IoU thresholds, with higher threshold having lower recall, as expected. Again, the LLM-based parser is better for obtaining the timestamps.
@@
Fig.  5 shows example detour localization (right) for our method. We see that our method is able to use the source video context and the user query to localize the detour window. For example, when a user asks “can I do this step with a steel plank saver”, our model correctly localizes the process of using a steel plank saver to put the salmon of the grill in the target video. See Supp. for more visualizations and comparisons.
@@ 5 Conclusion
We propose a novel task of finding detours for navigating instructional videos. Building on video and language modeling, we develop a weakly-supervised training dataset and a novel method to train a detours network. Our results show how existing similarity-based retrieval and localization methods are insufficient to address this problem. The dataset will be released to the community to support research in navigating instructional videos.
@@ 3 Detour dataset generation details
System: Help summarize the steps of this YouTube recipes whose narrations with timestamps are given. Timestamp is given in HH:MM:ss.
User: Given the narrations of a YouTube video, tell the recipe being made in this YouTube video and list down the steps and start and end timestamps in the video. Answer in this format: ‘Recipe: Name of the recipe and brief detail Step 1: [HH:MM:ss - HH:MM:ss] description of the step Step 2: [HH:MM:ss - HH:MM:ss] description of the step and so on’. Here are narrations with timestamps in HH:MM:ss format: $<$insert$>$
@@ 4.1 Generalizability to novel tasks
Method
Common Task MedR $\downarrow$
Novel Task MedR $\downarrow$
MedR $\downarrow$
@@
CLIP [61]
348
310
342
CLIP-Hitchhiker [7]
339
305
336
InternVideo [70]
315
296
313
CoVR [67]
464
485
473
Ours
29 35 30
Ours w/o hard-negatives
49
63
55
Ours w/ parser
76
88
81
Method
Common Task Mean R@1
Novel Task Mean R@1
Mean R@1
@@
2D-TAN [83]
8.9
8.2
8.6
VSLNet [81]
9.2
9.8
9.4
UMT [46]
9.6
9.3
9.4
STALE [57]
9.7
9.5
9.6
Ours
13.3 12.3 12.8
Ours w/ parser
12.0
11.3
11.6
@@
Method
$V_{s}$
$\mathcal{Q}$
R@5
R@10
R@50
MedR $\downarrow$
CLIP [61]
9.6
13.2
26.9
314
11.2
16.4
32.0
191
7.9
11.8
25.2
342
@@
11.3
17.7
33.2
186
8.4
12.3
25.6
336
InternVideo [70]
11.2
17.0
31.8
150
13.1
19.2
37.2
138
9.7
13.2
27.2
313
CoVR [67]
4.9
10.1
15.9
388
4.1
10.0
15.6
401
4.3
9.2
15.3
473
Ours
6.1
11.0
32.3
128
6.1
10.8
32.6
116
@@
Method
$V_{s}$
$\mathcal{Q}$
R@1, IoU=0.3
R@1, IoU=0.5
R@1, IoU=0.7
Mean R@1
2D-TAN [83]
8.9
3.2
0.9
5.5
10.0
3.8
1.2
8.0
10.3
4.2
1.5
8.6
VSLNet [81]
9.2
3.1
1.1
6.1
10.9
4.0
1.5
8.5
11.8
5.8
1.7
9.4
UMT [46]
9.7
3.5
1.2
6.5
11.2
5.4
1.7
8.7
12.0
6.1
1.6
9.4
STALE [57]
10.0
3.8
1.2
6.9
11.6
5.5
1.5
8.8
12.1
6.1
1.7
9.6
Ours
12.0
5.9
1.5
8.9
14.6
6.9
2.2
11.2

### v2

@@ 4.1 Detour video retrieval
First, we benchmark models on finding the correct detour video given the source video and the query, from amongst all videos in the test set (3,873 videos).
@@
These variants test whether individual embeddings (or simple combinations of them) are sufficient for detour retrieval.
@@
  • Distant Supervision [45] uses WikiHow as an external knowledge base to map steps in the video with keysteps. We replace the narrations with this keystep assignment for the detour dataset generation.
  • Multi-modal LLM [80] can be used to generate dense captions to replace the narrations. We use Vid2Seq [80] to first densely annotate the videos and convert the task to text-only video detours using those captions.
@@
Note that we train [45, 80, 69] on the same detour dataset as our model, whereas [72, 63, 7] are used in zero-shot setting for its good retrieval capabilities in instructional videos.
@@
Ablations. Similar to how we evaluate the baselines with different inputs, we compare the performance of our method (i.e. “with $V_{s},\mathcal{Q}$”) with ablations—“with $V_{s}$” and “with $\mathcal{Q}$”. Next, recall that we use a language model to generate timestamps for both the summaries and the detour timestamps (see Fig. 2 and Sec. 3.2), which can be noisy. Instead, for this ablation, we parse the summaries and queries and assign timestamps based on the similarity score between the narrations and the summaries. The language model does not handle timestamps for this “Ours w/ parser” baseline. Finally, we evaluate the retrieval performance without sampling negatives from the same task, i.e. no hard negatives in “Ours w/o hard negatives” baseline.
@@
Metrics. Following standard retrieval evaluation [55, 56, 75, 53], we report recall@$k$ for $k\in[5,10,50]$ and the median rank. Recall@$k$ ranges in $[0,1]$, higher the better. The median rank ranges between $1$ and the size of the retrieval candidate set ($3,873$), the lower the better.
@@
Results. Table 1 (left) reports the results with both inputs, i.e. “with $V_{s},\mathcal{Q}$” for all baselines on all metrics. We also report medR for different inputs in Table 2 (left). Our method outperforms all the baselines by a significant margin. Our performance gain is 7.9% w.r.t. InternVideo [72] at R@5 and the gap further increases to 35.2% for R@50. Across all the different input combinations, InternVideo [72] “with $\mathcal{Q}$” is the second best and attains a medR of 138 — much lower than our medR of 30 (Tab. 2).
@@
We attribute our large gain to three crucial factors: (a) thanks to its tokens and long sequence length in the LLM, our model is capable of long video understanding, unlike CoVR [69], which samples a few frames from the video; (b) appropriately using the source video context $V_{d}$, unlike CLIP [63], CLIP-Hitchhiker [7], and InternVideo [72], since our model is more than a standard text-to-video retrieval model; and (c) conditioning the retrieval on the query text. Furthermore, we find that using narrations for generating a weakly-supervised training set is better than using keysteps [45] or captions [80], which miss details typically present in narrations.
@@
Fig. 5 shows example detours inferred by our model. In all the successful cases, our retrieval model ranks the correct detour video at top-1. It is evident that detour video retrieval is more challenging than conventional text-to-video-retrieval because the detour query can have missing context, e.g. “how do I add garnish after pouring?” does not reveal the task in the video; the previous video context is needed for the correct retrieval. The failure cases (bottom row) show example errors in retrieval rank (left) and localization (right). The performance trend is similar for common and novel tasks of the dataset (see Supp.), reinforcing that training on a sufficiently large dataset with many recipes enables good generalization.
@@ 4.2 Detour window localization
Method $V_{s}$ $\mathcal{Q}$ MedR $\downarrow$
CLIP [63] 314
@@
Baselines. While there are no existing methods that use previous minutes-long viewing history along with the query to perform temporal localization, we use state-of-the-art localization methods as baselines and also strengthen them by providing source video context, as described below.
@@
  • 2D-TAN [86], VSLNet [84], UMT [47]: All these prior models aim to localize text in videos. To apply them for detour window localization, we train them to accept video context as input, namely using a visual mapper $f_{vm}$ similar to our approach—thus providing a late fusion of video $V_{s}$ context. These visual tokens are prepended to the text token, same as our token sequence (see Fig. 4). This enhancement enables us to evaluate “with $V_{s}$” and “with $V_{s}$, $\mathcal{Q}$” in addition to the standard “with $\mathcal{Q}$
@@
The baselines in [86, 84, 47] are trained on our same detour dataset for localization, whereas [58] is zero-shot.
@@
Results. Tab. 1 shows the results with greatest input context (“with $V_{s}$, $\mathcal{Q}$”), and Tab. 2 (right) compares against different input combinations. We outperform all the baselines and ablations by a clear margin. Our mean recall@1 is 3.2% higher than the second best performing method, STALE [58]. The same trend is true for all IoU thresholds, with higher thresholds having lower recall, as expected. Again, the LLM-based parser is better for obtaining the timestamps, and using narrations is better than an external knowledge base or captions.
@@
Fig. 5 also shows example detour localizations. We see that our method is able to use the source video context and the user query to localize the detour window. For example, when a user asks “can I do this step with a steel plank saver”, our model correctly localizes the use of a steel plank saver to put the salmon of the grill in the target video.
@@ 5 Conclusion
We propose a novel task of finding detours for navigating instructional videos. Building on video and language modeling, we develop a weakly-supervised training dataset and a novel method to train a detours network. Our results show how existing methods are insufficient to address this problem. The dataset will be released to the community to support research in navigating instructional videos.
Acknowledgements: UT Austin is supported in part by the IFML NSF AI Institute. KG is paid as a research scientist by Meta. We thank Suyog Jain, Austin Miller, Honey Manglani, and Robert Kuo for help with the data collection.
@@ 3 Detour dataset generation details
System: Help summarize the steps of this recipe whose narrations with timestamps are given. Timestamp is given in HH:MM:ss.
User: Given the narrations of a video, tell the recipe being made in this video and list down the steps and start and end timestamps in the video. Answer in this format: ‘Recipe: Name of the recipe and brief detail Step 1: [HH:MM:ss - HH:MM:ss] description of the step Step 2: [HH:MM:ss - HH:MM:ss] description of the step and so on’. Here are narrations with timestamps in HH:MM:ss format: $<$insert$>$
@@ 4.1 Generalizability to novel tasks
Method Common Task MedR $\downarrow$ Novel Task MedR $\downarrow$ MedR $\downarrow$
@@
CLIP [63] 348 310 342
CLIP-Hitchhiker [7] 339 305 336
InternVideo [72] 315 296 313
DistantSup. [45] 320 350 329
MLLM [80] 127 155 139
CoVR [69] 464 485 473
Ours 29 35 30
Ours w/o hard-negatives 49 63 55
Ours w/ parser 76 88 81
Method Common Task Mean R@1 Novel Task Mean R@1 Mean R@1
@@
2D-TAN [86] 8.9 8.2 8.6
VSLNet [84] 9.2 9.8 9.4
UMT [47] 9.6 9.3 9.4
DistantSup. [45] 8.8 7.9 8.3
MLLM [80] 9.7 10.8 10.2
STALE [58] 9.7 9.5 9.6
Ours 13.3 12.3 12.8
Ours w/ parser 12.0 11.3 11.6
@@
Method $V_{s}$ $\mathcal{Q}$ R@5 R@10 R@50 MedR $\downarrow$
CLIP [63] 9.6 13.2 26.9 314
11.2 16.4 32.0 191
7.9 11.8 25.2 342
@@
11.3 17.7 33.2 186
8.4 12.3 25.6 336
InternVideo [72] 11.2 17.0 31.8 150
13.1 19.2 37.2 138
9.7 13.2 27.2 313
DistantSup. [45] 4.9 10.2 15.9 384
8.0 12.0 25.4 370
8.4 12.6 25.1 329
MLLM [80] 11.3 17.8 32.3 189
11.4 16.8 31.4 158
5.9 10.5 32.1 139
CoVR [69] 4.9 10.1 15.9 388
4.1 10.0 15.6 401
4.3 9.2 15.3 473
Ours 6.1 11.0 32.3 128
6.1 10.8 32.6 116
@@
Method $V_{s}$ $\mathcal{Q}$ R@1, IoU=0.3 R@1, IoU=0.5 R@1, IoU=0.7 Mean R@1
2D-TAN [86] 8.9 3.2 0.9 5.5
10.0 3.8 1.2 8.0
10.3 4.2 1.5 8.6
VSLNet [84] 9.2 3.1 1.1 6.1
10.9 4.0 1.5 8.5
11.8 5.8 1.7 9.4
UMT [47] 9.7 3.5 1.2 6.5
11.2 5.4 1.7 8.7
12.0 6.1 1.6 9.4
DistantSup. [45] 9.8 3.7 1.2 7.6
10.0 3.8 1.2 7.9
10.6 4.0 1.5 8.3
MLLM [80] 12.2 6.0 1.5 9.1
12.3 6.2 1.7 9.7
12.7 6.5 1.8 10.2
STALE [58] 10.0 3.8 1.2 6.9
11.6 5.5 1.5 8.8
12.1 6.1 1.7 9.6
Ours 12.0 5.9 1.5 8.9
14.6 6.9 2.2 11.2